# Magnon transmission across $\nu=1|-1|1$ mono-layer graphene junction as a probe of electronic structure

### v1

@@ III Hamiltonian and HF approximations with junctions
$\mathcal{H}_{HF}=\frac{\ell}{L_{y}}\sum_{k_{1},k_{2}}\sum_{i,j}\big{[}\sum_{m,n}(V_{ijmn}(k_{1}-k_{2})-$
$V_{inmj}(k_{1}-k_{2}))\Delta_{mn}(k_{2})+V_{H}(k_{1}-k_{2})\Delta_{jj}(k_{2})$
$-V_{F}(k_{1}-k_{2})\Delta_{ji}(k_{2})\big{]}c_{k_{1}i}^{\dagger}c_{k_{1}j}-\sum_{k}V_{bg}(k)\vec{c}_{k}^{\hskip 2.84544pt\dagger}\vec{c}_{k}$
$-\sum_{k}E_{z}\vec{c}_{k}^{\hskip 2.84544pt\dagger}\sigma_{z}\vec{c}_{k}-\sum_{k}E_{V}\vec{c}_{k}^{\hskip 2.84544pt\dagger}\tau_{z}\vec{c}_{k}\hskip 65.44142pt$ (10)
@@
Figure 3: HF energies (top panel), spin structure (middle panel), and valley structure (bottom panel) of each self-consistent HF level for ${B_{\perp}}=4{B_{\perp}}^{0}$. The Hamiltonian parameters are ${\tilde{W}}^{0}=20,~{}E_{C}^{0}=30,~{}E_{Z}^{0}=0.5,~{}E_{V}=5.0,~{}u_{z}^{0}=4,~{}u_{\perp}^{0}=-3$. The spin and valley structure of each HF level is shown by plotting the vectors $\vec{\sigma}=(\langle\sigma_{x}\rangle,\langle\sigma_{z}\rangle)$(for spin) and $\vec{\tau}=(\langle\tau_{x}\rangle,\langle\tau_{z}\rangle)$(for valley). $\langle\sigma_{z}\rangle=+1,-1$ is identified with $\uparrow$,$\downarrow$ spin and $\langle\tau_{z}\rangle=+1,-1$ with the valleys $K$,$K^{\prime}$ respectively. The direction of the arrows of spin and valley represent the orientation of the averages of $\vec{\tau},~{}\vec{\sigma}$ in the $xz$ plane of each internal space. The colors of the arrows are the same as that of the corresponding energy levels. For the parameters chosen, the bulk $\nu=1$ ordering is $E_{K^{\prime}\uparrow}<E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$ and the self-consistent ground state prefers many-body spin rotation, whereas the valley indices of each level flip discontinuously across the interfaces.
@@
We will present some HF results here to explicitly show how the bulk ordering in $\nu=1$, which is dictated by the ratio of $u_{z}^{0}$ and $u_{\perp}^{0}$ and ${B_{\perp}}$ (for fixed $E_{V},E_{Z}^{0}$), determines the self-consistent HF state of the junction. In Landau gauge, the momentum $k$ along the periodic direction $y$ is related to the guiding center position $X_{k}=k\ell^{2}$ along the $x$-direction. The figures we present in what follows show the HF single-particle energies and the spin-valley directions of each HF state as functions of the guiding center position $X_{k}$. Since we have used the $U(1)_{s}\otimes U(1)_{v}$ symmetries to make $\Delta$ real, the averages of $\vec{\tau},~{}\vec{\sigma}$ lie in the $xz$ plane in each internal space. We will simply present the directions as arrows with an $\uparrow$ representing $K$ in the case of valley, and $\uparrow$ representing $\langle\sigma_{z}\rangle=1$ in the case of spin.
@@
Figure 4: HF energies (top panel), spin structure (middle panel), and valley structure (bottom panel) of each self-consistent HF level for ${B_{\perp}}=4{B_{\perp}}^{0}$. The Hamiltonian parameters are ${\tilde{W}}^{0}=20,~{}E_{C}^{0}=30,~{}E_{Z}^{0}=0.5,~{}E_{V}=5.0,~{}u_{z}^{0}=2,~{}u_{\perp}^{0}=-3$. We have used the same convention to plot the spin and valley structure here as in Fig. (3). In this case, the bulk ordering of the energy levels $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, is different from the case in Fig. (3). Here, we find self-consistently that the ground state prefers valley rotation and the spin flips discontinuously across the interfaces.
@@
First, consider the case where $u_{z}^{0}>|u_{\perp}^{0}|$, shown in Fig. (3). We have chosen the parameters ${B_{\perp}}/{B_{\perp}}^{0}=4$ with the other parameters identical to those of Fig. (1). The ordering of the HF levels deep in the bulk of $\nu=1$ is $E_{K^{\prime}\uparrow}<E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, whereas the filled state for bulk $\nu=-1$ is ${K\uparrow}$. As can be seen from the directions of the spin and valley for each single-particle state, the system prefers to spontaneously break the $U(1)_{s}$ symmetry at each interface, rotating the spins continuously. However, the valley degree of freedom remains polarized either at either $K$ or $K^{\prime}$, keeping the $U(1)_{v}$ symmetry intact. A level crossing occurs at the interface between the two lowest levels, discontinuously exchanging their valley polarizations.
@@
Next, let us consider the case $u_{z}^{0}<|u_{\perp}^{0}|$, shown in Fig. (4). For ${B_{\perp}}/{B_{\perp}}^{0}=4$ with the other parameters being chosen as in Fig. (2), the ordering of the HF levels in the $\nu=1$ bulk is $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, whereas the filled state for bulk $\nu=-1$ is ${K\uparrow}$. Now a spontaneous breaking of the $U(1)_{v}$ symmetry occurs, and there is a continuous rotation of the valley polarization of each single-particle state. However, the spin directions remain frozen at either $\uparrow$ or $\downarrow$, with an abrupt change occuring at the interfaces via a level crossing between the two lowest levels.
@@
The above two examples show the importance of the ratio $u_{z}^{0}/|u_{\perp}^{0}|$ to the structure of the interface. We will present many more examples in Section V and show how the structure of the interfaces in turn affects magnon transmission across the junction. However, we will first examine the collective excitations of the system and review the magnon scattering formalism [47] using the time-dependent Hartree-Fock (TDHF) method.
@@ IV.1 Bulk collective excitations of $\nu=1$
Figure 9: Occupied single-particle HF energies for bulk $\nu=1$ as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=1$, $u_{z}^{0}=4$ and $u_{\perp}^{0}=-3$. The parameters are chosen such that it reflects the case when $u_{z}^{0}>|u_{\perp}^{0}|>E_{V}>E_{Z}^{0}$. As we can see, for this case the ordering of the occupied state is $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}$ which does not change with increasing ${B_{\perp}}$.
@@
Figure 12: Occupied single-particle HF energies for the bulk $\nu=1$ state as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=1$, $u_{z}^{0}=2$ and $u_{\perp}^{0}=-3$. We have chosen the parameters such that it reflects the case when $|u_{\perp}^{0}|>u_{z}^{0}>E_{V}>E_{Z}^{0}$. In this case we find the bulk ordering is $E_{K{\downarrow}}<E_{K^{\prime}{\uparrow}}<E_{K{\uparrow}}$, which does not change with ${B_{\perp}}$.
@@
are the self-energy contributions which arise from integrating out the asymptotic regions. With $\beta$ running over all the indices $(m,n)$ of the bulk collective modes. $\mathbb{K}_{X_{i},m,n}^{X_{i^{\prime}},m^{\prime},n^{\prime}}(q_{y})$ is given in Eq. (20) with $X_{i}=k\ell^{2}$ and $X_{i^{\prime}}=k^{\prime}\ell^{2}$. The inhomogeneous term is given by
@@
Figure 15: Occupied Hartree-Fock single-particle energies for the bulk $\nu=1$ as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=0.1$, $u_{z}^{0}=4$ and $u_{\perp}^{0}=-3$. The parameters satisfy $u_{z}^{0}>|u_{\perp}^{0}|>E_{Z}^{0}>E_{V}$. In this case the ordering of the occupied levels is $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}$ which remains unchanged with increasing ${B_{\perp}}$.
@@
Figure 18: Occupied Hartree-Fock single-particle energies for the bulk $\nu=1$ as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=0.1$, $u_{z}^{0}=2$ and $u_{\perp}^{0}=-3$. The parameters satisfy $|u_{\perp}^{0}|>u_{z}^{0}>E_{Z}^{0}>E_{V}$. For this case the ordering of the occupied states is $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}$ which does not change with increasing ${B_{\perp}}$.
@@
In view of the large number of parameters in the problem, $E_{V},~{}B_{\perp},~{}u^{0}_{z},~{}u^{0}_{\perp},{\tilde{W}}^{0}$, we need to organize the results. The case $E_{V}=0$ was already addressed in Ref. [47], and hence we will always take $E_{V}>0$ in what follows. We will focus solely on the regime of short-range interactions where $u^{0}_{z}>0,~{}u^{0}_{\perp}<0$, believed to be realized in graphene. Here there are two major cases, (i) $u_{z}>|u_{\perp}|$ and (ii) $u_{z}<|u_{\perp}|$. Our second organizing principle is to start with small ${B_{\perp}}$ and go towards large ${B_{\perp}}$. At small $B_{\perp}$, for samples $E_{V}$ will typically be larger than the short-range couplings (due to the dependence of $u_{z},~{}u_{\perp}$ on $B_{\perp}$), whereas at larger $B_{\perp}$ the short-range couplings may be of the same order or even larger than $E_{V}$. The relative magnitudes of the short-range couplings vs $E_{V}$ have profound effects on the self-consistent structure of the interfaces, and thus on the magnon transmission amplitudes across the system. One of the primary motivations of this work is to use such as study to constrain the ratio $u_{z}/|u_{\perp}|$.
In view of these considerations, this section is organized as follows: In Subsection V.1, we examine the case $E_{V}>u^{0}_{z},|u^{0}_{\perp}|$ over a range of $B_{\perp}$. Results for both $u_{z}>|u_{\perp}|$ and $u_{z}<|u_{\perp}|$ will be presented. We find that for $u_{z}>|u_{\perp}|$ there is no qualitative change in the magnon transmission amplitudes as $B_{\perp}$ increases, while for $u_{z}<|u_{\perp}|$ the magnon transmission amplitude at high energy changes dramatically as $B_{\perp}$ increases. Interestingly, in this latter case, a spin-valley magnon is also excited. Next, in Subsection V.2, we examine the case of intermediate $E_{V}$, with the same subcases as in Subsection V.1. Finally, in Subsection V.3, we consider the case when $E_{V}$ is the smallest energy scale in the problem, with the inequalities $|u_{\alpha}^{0}|>E_{Z}^{0}>E_{V}$.
@@ V.2 Magnon transmission at intermediate $E_{V}$
First, we consider the case where $u_{z}^{0}>|u_{xy}^{0}|$. We illustrate this case with the parameters $u_{z}^{0}=4,~{}u_{\perp}^{0}=-3,~{}E_{V}=1,~{}E_{Z}=0.5$, for which the one-body energies are shown as a function of $B_{\perp}$ in Fig. 9. As can be seen, we always have $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}<E_{K^{\prime}{\downarrow}}$ for $\nu=1$. The occupied state for $\nu=-1$ is ${|K\uparrow\rangle}$.
The self-consistent HF solutions for ${B_{\perp}}={B_{\perp}}^{0},2{B_{\perp}}^{0},4{B_{\perp}}^{0}$ are shown in Fig.  10. As in the case of large $E_{V}$, we find that for $u_{z}>|u_{\perp}|$ the system prefers to undergo spin rotations at the interfaces, leaving the valley quantum number conserved. It is therefore not surprising that the transmission amplitudes, shown in Fig.  11, are very similar to those at large $E_{V}$ (Fig. 6). There are sharp peaks at low energies, which we believe represent coupling to the cavity collective modes in the $\nu=-1$ region mediated by the interface collective modes. At high energies the transmission drops to zero.
Next, we consider the case where $u_{z}^{0}<|u_{\perp}^{0}|$. We have illustrated this case with the following choice of parameters: $u_{z}^{0}=2,~{}u_{\perp}^{0}=-3,~{}E_{V}=1,~{}E_{Z}=0.5$. As shown in Fig. 12, the ordering of the one-body energies does not change as $B_{\perp}$ increases.

### v2

@@ III Hamiltonian and HF approximations with junctions
$\mathcal{H}_{HF}=\frac{\ell}{L_{y}}\sum_{k_{1},k_{2}}\sum_{i,j}\big{[}\sum_{m,n}\big{(}V_{ijmn}(k_{1}-k_{2})~{}-$
$V_{inmj}(k_{1}-k_{2})\big{)}\Delta_{mn}(k_{2})c_{k_{1}i}^{\dagger}c_{k_{1}j}~{}+$
$V_{H}(k_{1}-k_{2})\Delta_{jj}(k_{2})c_{k_{1}i}^{\dagger}c_{k_{1}i}~{}-$
$V_{F}(k_{1}-k_{2})\Delta_{ji}(k_{2})c_{k_{1}i}^{\dagger}c_{k_{1}j}\big{]}~{}-$
$\sum_{k}V_{bg}(k)\vec{c}_{k}^{\hskip 2.84544pt\dagger}\vec{c}_{k}-\sum_{k}E_{Z}\vec{c}_{k}^{\hskip 2.84544pt\dagger}\sigma_{z}\vec{c}_{k}-\sum_{k}E_{V}\vec{c}_{k}^{\hskip 2.84544pt\dagger}\tau_{z}\vec{c}_{k}$ (10)
@@
Figure 3: HF energies (top panel), spin structure (middle panel), and valley structure (bottom panel) of each self-consistent HF level for ${B_{\perp}}=4B_{\perp}^{0}$. The Hamiltonian parameters are ${\tilde{W}}^{0}=20,~{}E_{C}^{0}=30,~{}E_{Z}^{0}=0.5,~{}E_{V}=5.0,~{}u_{z}^{0}=4,~{}u_{xy}^{0}=-3$. The spin and valley structure of each HF level is shown by plotting the vectors $\vec{\sigma}=(\langle\sigma_{x}\rangle,\langle\sigma_{z}\rangle)$(thin arrows for spin) and $\vec{\tau}=(\langle\tau_{x}\rangle,\langle\tau_{z}\rangle)$(thick arrows for valley). $\langle\sigma_{z}\rangle=+1,-1$ is identified with $\uparrow$,$\downarrow$ spin and $\langle\tau_{z}\rangle=+1,-1$ with the valleys $K$,$K^{\prime}$ respectively. The direction of the arrows of spin and valley represent the orientation of the averages of $\vec{\tau},~{}\vec{\sigma}$ in the $xz$ plane of each internal space. The colors of the arrows are the same as that of the corresponding energy levels. For the parameters chosen, the bulk $\nu=1$ ordering is $E_{K^{\prime}\uparrow}<E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$ and the self-consistent ground state prefers many-body spin rotation, whereas the valley indices of each level flip discontinuously across the interfaces.
@@
We will present some HF results here to explicitly show how the bulk ordering in $\nu=1$, which is dictated by the ratio of $u_{z}^{0}$ and $u_{xy}^{0}$ and ${B_{\perp}}$ (for fixed $E_{V},E_{Z}^{0}$), determines the self-consistent HF state of the junction. In Landau gauge, the momentum $k$ along the periodic direction $y$ is related to the guiding center position $X_{k}=k\ell^{2}$ along the $x$-direction. The figures we present in what follows show the HF single-particle energies and the spin-valley directions of each HF state as functions of the guiding center position $X_{k}$. Since we have used the $U(1)_{s}\otimes U(1)_{v}$ symmetries to make $\Delta$ real, the averages of $\vec{\tau},~{}\vec{\sigma}$ lie in the $xz$ plane in each internal space. We will simply present the directions as arrows with an $\uparrow$ representing $K$ in the case of valley, and $\uparrow$ representing $\langle\sigma_{z}\rangle=1$ in the case of spin.
@@
Figure 4: HF energies (top panel), spin structure (middle panel), and valley structure (bottom panel) of each self-consistent HF level for ${B_{\perp}}=4B_{\perp}^{0}$. The Hamiltonian parameters are ${\tilde{W}}^{0}=20,~{}E_{C}^{0}=30,~{}E_{Z}^{0}=0.5,~{}E_{V}=5.0,~{}u_{z}^{0}=2,~{}u_{xy}^{0}=-3$. We have used the same convention to plot the spin and valley structure here as in Fig. (3), with thin arrows for spin and thick arrows for valley. In this case, the bulk ordering of the energy levels $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, is different from the case in Fig. (3). Here, we find self-consistently that the ground state prefers valley rotation and the spin flips discontinuously across the interfaces.
@@
First, consider the case where $u_{z}^{0}>|u_{xy}^{0}|$, shown in Fig. (3). We have chosen the parameters ${B_{\perp}}/B_{\perp}^{0}=4$ with the other parameters identical to those of Fig. (1). The ordering of the HF levels deep in the bulk of $\nu=1$ is $E_{K^{\prime}\uparrow}<E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, whereas the filled state for bulk $\nu=-1$ is ${K\uparrow}$. As can be seen from the directions of the spin and valley for each single-particle state, the system prefers to spontaneously break the $U(1)_{s}$ symmetry at each interface, rotating the spins continuously. However, the valley degree of freedom remains polarized either at $K$ or $K^{\prime}$, keeping the $U(1)_{v}$ symmetry intact. A level crossing occurs at the interface between the two lowest levels, discontinuously exchanging their valley polarizations.
@@
Next, let us consider the case $u_{z}^{0}<|u_{xy}^{0}|$, shown in Fig. (4). For ${B_{\perp}}/B_{\perp}^{0}=4$ with the other parameters being chosen as in Fig. (2), the ordering of the HF levels in the $\nu=1$ bulk is $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}<E_{K^{\prime}\downarrow}$, whereas the filled state for bulk $\nu=-1$ is ${K\uparrow}$. Now a spontaneous breaking of the $U(1)_{v}$ symmetry occurs, and there is a continuous rotation of the valley polarization of each single-particle state. However, the spin directions remain frozen at either $\uparrow$ or $\downarrow$, with an abrupt change occuring at the interfaces via a level crossing between the two lowest levels.
@@
The above two examples show the importance of the ratio $u_{z}^{0}/|u_{xy}^{0}|$ to the structure of the interface. We will present many more examples in Section V and show how the structure of the interfaces in turn affects magnon transmission across the junction. However, we will first examine the collective excitations of the system and review the magnon scattering formalism [47] using the time-dependent Hartree-Fock (TDHF) method.
@@ IV.1 Bulk collective excitations of $\nu=1$
Figure 9: Occupied single-particle HF energies for bulk $\nu=1$ as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=1$, $u_{z}^{0}=4$ and $u_{xy}^{0}=-3$. The parameters are chosen such that it reflects the case when $u_{z}^{0}>|u_{xy}^{0}|>E_{V}>E_{Z}^{0}$. As we can see, for this case the ordering of the occupied state is $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}$ which does not change with increasing ${B_{\perp}}$.
@@
Figure 12: Occupied single-particle HF energies for the bulk $\nu=1$ state as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=1$, $u_{z}^{0}=2$ and $u_{xy}^{0}=-3$. We have chosen the parameters such that it reflects the case when $|u_{xy}^{0}|>u_{z}^{0}>E_{V}>E_{Z}^{0}$. In this case we find the bulk ordering is $E_{K{\downarrow}}<E_{K^{\prime}{\uparrow}}<E_{K{\uparrow}}$, which does not change with ${B_{\perp}}$.
@@
are the self-energy contributions which arise from integrating out the asymptotic regions. With $\beta$ running over all the indices $(m,n)$ of the bulk collective modes. $\mathbb{K}_{X_{i},m,n}^{X_{i^{\prime}},m^{\prime},n^{\prime}}(q_{y})$ is given in Eq. (20) with $X_{i}=k\ell^{2}$ and $X_{i^{\prime}}=k^{\prime}\ell^{2}$. The inhomogeneous source term is given by
@@
Figure 15: Occupied Hartree-Fock single-particle energies for the bulk $\nu=1$ as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=0.1$, $u_{z}^{0}=4$ and $u_{xy}^{0}=-3$. The parameters satisfy $u_{z}^{0}>|u_{xy}^{0}|>E_{Z}^{0}>E_{V}$. In this case the ordering of the occupied levels is $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}$ which remains unchanged with increasing ${B_{\perp}}$.
@@
Figure 18: Occupied Hartree-Fock single-particle energies for the bulk $\nu=1$ as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$, $E_{Z}^{0}=0.5$, $E_{V}=0.1$, $u_{z}^{0}=2$ and $u_{xy}^{0}=-3$. The parameters satisfy $|u_{xy}^{0}|>u_{z}^{0}>E_{Z}^{0}>E_{V}$. For this case the ordering of the occupied states is $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}$ which does not change with increasing ${B_{\perp}}$.
@@
In view of the large number of parameters in the problem, $E_{V},~{}B_{\perp},~{}u^{0}_{z},~{}u_{xy}^{0},{\tilde{W}}^{0}$, we need to organize the results. The case $E_{V}=0$ was already addressed in Ref. [47], and hence we will always take $E_{V}>0$ in what follows. We will focus solely on the regime of short-range interactions where $u^{0}_{z}>0,~{}u_{xy}^{0}<0$, believed to be realized in graphene. Here there are two major cases, (i) $u_{z}>|u_{xy}|$ and (ii) $u_{z}<|u_{xy}|$. Our second organizing principle is to start with small ${B_{\perp}}$ and go towards large ${B_{\perp}}$. At small $B_{\perp}$, for samples $E_{V}$ will typically be larger than the short-range couplings (due to the dependence of $u_{z},~{}u_{xy}$ on $B_{\perp}$), whereas at larger $B_{\perp}$ the short-range couplings may be of the same order or even larger than $E_{V}$. The relative magnitudes of the short-range couplings vs $E_{V}$ have profound effects on the self-consistent structure of the interfaces, and thus on the magnon transmission amplitudes across the system. One of the primary motivations of this work is to use such as study to constrain the ratio $u_{z}/|u_{xy}|$.
In view of these considerations, this section is organized as follows: In Subsection V.1, we examine the case $E_{V}>u^{0}_{z},|u_{xy}^{0}|$ over a range of $B_{\perp}$. Results for both $u_{z}>|u_{xy}|$ and $u_{z}<|u_{xy}|$ will be presented. We find that for $u_{z}>|u_{xy}|$ there is no qualitative change in the magnon transmission amplitudes as $B_{\perp}$ increases, while for $u_{z}<|u_{xy}|$ the magnon transmission amplitude at high energy changes dramatically as $B_{\perp}$ increases. Interestingly, in this latter case, a spin-valley magnon is also excited. Next, in Subsection V.2, we examine the case of intermediate $E_{V}$, with the same subcases as in Subsection V.1. Finally, in Subsection V.3, we consider the case when $E_{V}$ is the smallest energy scale in the problem, with the inequalities $|u_{\alpha}^{0}|>E_{Z}^{0}>E_{V}$.
@@ V.2 Magnon transmission at intermediate $E_{V}$
First, we consider the case where $u_{z}^{0}>|u_{xy}^{0}|$. We illustrate this case with the parameters $u_{z}^{0}=4,~{}u_{xy}^{0}=-3,~{}E_{V}=1,~{}E_{Z}=0.5$, for which the one-body energies are shown as a function of $B_{\perp}$ in Fig. 9. As can be seen, we always have $E_{K^{\prime}{\uparrow}}<E_{K{\downarrow}}<E_{K{\uparrow}}<E_{K^{\prime}{\downarrow}}$ for $\nu=1$. The occupied state for $\nu=-1$ is ${|K\uparrow\rangle}$.
The self-consistent HF solutions for ${B_{\perp}}=B_{\perp}^{0},2B_{\perp}^{0},4B_{\perp}^{0}$ are shown in Fig.  10. As in the case of large $E_{V}$, we find that for $u_{z}>|u_{xy}|$ the system prefers to undergo spin rotations at the interfaces, leaving the valley quantum number conserved. It is therefore not surprising that the transmission amplitudes, shown in Fig.  11, are very similar to those at large $E_{V}$ (Fig. 6). There are sharp peaks at low energies, which we believe represent coupling to the cavity collective modes in the $\nu=-1$ region mediated by the interface collective modes. At high energies the transmission drops to zero.
Next, we consider the case where $u_{z}^{0}<|u_{xy}^{0}|$. We have illustrated this case with the following choice of parameters: $u_{z}^{0}=2,~{}u_{xy}^{0}=-3,~{}E_{V}=1,~{}E_{Z}=0.5$. As shown in Fig. 12, the ordering of the one-body energies does not change as $B_{\perp}$ increases.